# Causality for Earth Science - A Review on Time-series and Spatiotemporal Causality Methods

### v1

@@ 3.3 Common Approaches
Causal discovery methods can be categorized into following three major categories or approaches:
@@
One of the most common approaches for performing causal discovery is the constraint-based approach which relies on testing the conditional independencies between the variables to identify the causal edges. In the constraint-based category, one of the oldest approaches is the Peter-Clark (PC) algorithm proposed by (Spirtes et al., 2000). It assumes faithfulness which implies that all the interdependencies in a DAG need to be under the d-separation criterion. Another well-known constraint-based approach is the Fast Causal Inference (FCI) algorithm (Spirtes et al., 2000) which differs from the PC algorithm as it violates causal sufficiency. That is, FCI assumes that some common causes may be unmeasured in real-world data.
@@
The score-based approach is another common approach for causal discovery. Score-based approaches try to find the DAG with the maximum likelihood given the data (Triantafillou and Tsamardinos, 2016). They search over the space of all possible DAGs aiming to maximize a score function that depicts how well the graph fits the data. As a score function, the Bayesian Information Criterion (BIC) and other similar approaches are used to score the produced graphs and output the DAG with the best score (fits data the best). The score function can be modified to optimize or reduce the search space of all possible DAGs. In terms of performance, score-based algorithms can be cost ineffective as they need to score every possible DAG (Nogueira et al., 2021). When the number of variables increases even slightly, the number of possible graphs can grow super-exponentially. One of the famous score-based methods is the Greedy Equivalence Search (GES) algorithm (Chickering, 2002) that initially starts with an empty graph, and keeps on adding currently needed edges, and then deletes unnecessary edges as per the score function.
@@
Functional Causal Model (FCM) or also called SCM based approaches express the causal relationships in a specific functional form (a set of equations) where the variables are a function ($y=f(pa_{y})+e$) of their direct causes (causal parents) and an independent noise term $e$. These approaches rely heavily on the independent noise for the recovery and estimation of the causal relations. The FCM-based methods try to distinguish different DAGs in the same equivalence class by imposing additional assumptions on the data distributions and/or function classes. A wide range of FCM-based approaches exists for both temporal and non-temporal data that tries to handle both linear and non-linear causal relations using methods such as independent component analysis (ICA) (Stone, 2004) and additive noise models (ANMs) (Hoyer et al., 2008) respectively.
@@ 3.4 Time-Series Causal Discovery
Timeseries causal discovery refers to the process of identifying causal relationships between variables in a timeseries data setting. Causal discovery in timeseries data involves dealing with the specific challenges of temporal dependencies and potential feedback loops. It has been widely used in econometrics, neuroscience, and climate science (Spirtes et al., 2000; Zhou et al., 2014; Van Nes et al., 2015; Ghysels et al., 2016). Identifying time lags is an important aspect of causal discovery in timeseries data. It involves determining the delay or time difference between the cause and its effect. Time lags are crucial for understanding the dynamics of the causal relationship and for making accurate predictions in timeseries forecasting tasks. To identify time lags in causal relationships, researchers may use methods like cross-correlation analysis, autocorrelation, or timeseries cross-validation techniques (Spirtes et al., 2000). These methods help measure the strength of the association between variables at different time lags and assist in determining the optimal lag that maximizes the predictive performance or the strength of the causal relationship. We explain some of the widely used causal discovery methods below, along with their variants, underlying causal assumptions and limitations. A tabulated summary of these time-series causal discovery methods is given in Table 5.
@@ 3.4.1 Granger Causality, V-Granger
Granger causality is a statistical hypothesis test used to determine whether one time series can predict another Granger (1969). It is named after Clive Granger, a Nobel laureate economist who developed the concept. Granger causality is widely used in time series analysis to investigate the causal relationships between variables. The basic idea behind Granger causality is that if a time series X ”Granger-causes” another time series Y, then past values of X contain information that helps predict future values of Y beyond what can be predicted using past values of Y alone. It is important to note that Granger causality does not establish a true cause-and-effect relationship in the sense of a controlled experiment. It only identifies whether one time series provides useful predictive information for another time series. Nevertheless, Granger causality is commonly used in various fields, including economics, finance, engineering, and neuroscience, to explore relationships between time-varying data. However, it has limitations and assumptions. For instance, the choice of time lags can influence the results. GC assumes causal sufficiency condition, stationary and linear relationship and does not predict the direction of causal relation. The method is data hungry and requires sufficiently large sample size. Granger causality can be a useful exploratory tool, and its results should be interpreted carefully in the context of the specific data and research question at hand.
With the Granger Causality widely used in various fields to understand the causal relationship between time series, the presence of unobserved confounders poses a fundamental problem in observational studies, especially for non-linear Granger causality. V-Granger (Meng, 2019), overcomes the causal sufficiency limitation by estimating Granger causality in the presence of unobserved confounders using a variational autoencoder (VAE) to estimate the intractable posterior distribution and a recurrent neural network to model the temporal relationship in the data. Through experiments on synthetic and semi-synthetic datasets, as well as two real datasets: Butter and Temperature, the authors demonstrated how V-Granger outperforms the traditional Granger test in some cases. However, the method is not suitable for datasets with high noise levels.
@@ 3.4.2 PCMCI, PCMCI+, LPCMCI
PCMCI1 (Runge et al., 2019a) is a method for efficient estimation of causal graphs from high-dimensional time series datasets, which can be used for robust forecasting and the estimation and prediction of direct, total, and mediated effects. A baseline of PCMCI was first introduced by (Runge et al., 2015) for identifying causal gateways and mediators in complex spatiotemporal systems. The method combines the PC stable causal discovery algorithm with momentary conditional independence (MCI) to estimate causal networks from large-scale time series datasets in a two-step approach. The first step is to find lagged parents of all individual nodes using the PC stable method. The second step is to test the momentary conditional independence. As a result of the second step, if two nodes are independent, then there exists no cause-effect relationship between them. The outcome of this two-step method includes a causal adjacency matrix, corresponding lag values and strength of identified relationships given by p-values. PCMCI supports multiple linear and non-parametric conditional independence tests depending on the data distribution. For instance, the partial correlation or $ParCorr()$ test assumes linear dependencies and Gaussian noise in the data. There are multiple variants of PCMCI available for different downstream tasks. For instance, PCMCI+ can be used for discovering contemporaneous and lagged causal relations in autocorrelated nonlinear time series datasets. For details on PCMCI’s implementation, independence tests available and different variants of PCMCI, please refer to their Github [^1] https://github.com/jakobrunge/tigramite repository.
Similarly, Latent PCMCI or LPCMCI1 (Gerhardus and Runge, 2020) is a constraint-based approach which is a variant of the PCMCI algorithm for causal discovery from large-scale time series data in the presence of hidden confounders. This method is specialized to learn lag-specific causal relationships for linear and nonlinear, lagged and contemporaneous causal discovery from observational time series data.
@@ 3.4.3 LinGAM, VARLiNGAM, TS-LiNGAM
LinGAM [^2] https://lingam.readthedocs.io/en/latest/index.html (Linear Non-Gaussian Acyclic Model) (Shimizu et al., 2006) is a framework used for causal discovery and estimation of causal effects from observational data. It focuses on identifying linear causal relationships among variables while considering non-Gaussian noise distributions. LinGAM assumes acyclic causal relationships among variables, meaning it aims to identify a directed acyclic graph (DAG) that represents the underlying causal structure. VARLiNGAM Hyvärinen et al. (2010) extends the foundational LiNGAM model to encompass scenarios involving time series. This augmentation involves integrating the core LiNGAM model with the conventional vector autoregressive models (VAR). Consequently, it facilitates the exploration of causal relationships encompassing both delayed and concurrent (instantaneous) effects. This stands in contrast to the traditional VAR framework, which solely examines causal relations with a time delay. VARLiNGAM holds similar causal assumptions as the base LinGAM model, such as, data linearity, causal sufficiency with no hidden confounders, acyclicity of contemporaneous causal relations and non-gaussian continuous error variables.
TS-LiNGAM (Hyvärinen et al., 2008) is a generalization of the LiNGAM algorithm which uses a non-Gaussianity assumption to estimate both instantaneous and lagged causal edges. The model is a combination of autoregressive models and structural equation modeling (SEM). It assumes the external noises to be mutually independent, temporally uncorrelated, and to be non-Gaussian. TS-LiNGAM combines classic least-squares estimation of an autoregressive (AR) model with LiNGAM estimation. It does so in four steps: i) Estimation of a classic autoregressive model, ii) Residuals computation, iii) Performing LiNGAM analysis on the residuals, and iv) Computation of the causal effect estimates. The method has been evaluated on real datasets such as financial (stock) data and magnetoencephalographic data i.e measurements of the electric activity in the brain.
@@ 3.4.4 TiMiNo
TiMINo (Peters et al., 2013) approach provides a new method at that time for causal inference on time series data that overcomes some of the methodological issues of previous methods that could not capture nonlinear and instantaneous effects. Specifically, TiMINo extends Structural Equation Models (SEMs) to the time series setting, and uses a restricted function class to provide general identifiability results including lagged and instantaneous effects that can be nonlinear and unfaithful, and non-instantaneous feedback between the time series. Given a sample of finite and continuous-distributed multivariate time series data, and output the summary time graph if the input data satisfies model assumptions. Otherwise, if the input fails the assumptions, TiMINo causality remains mostly undecided instead of drawing wrong causal conclusions. The model is defined as $X_{t}^{i}=f_{i}((\mathbf{PA_{p}^{i}})_{t-p},\ldots,(\mathbf{PA_{1}^{i}})_{t-1},(\mathbf{PA_{0}^{i}})_{t},N_{t}^{i})$, where $p$ is the time lag for independence test, $(\mathbf{PA_{1}^{i}})$ is the set of parent nodes of node $i$, $X_{t}$ is a time-series data with finite and continuous distribution, and $N_{t}^{i}$ is a noise vector jointly independent over i and t. The approach assumes that each time series is a function of all direct causes and some noise variable, and restricts the function class to provide general identifiability results. Potential weaknesses of TiMINo are discussed including the possibility of fitting a model in the wrong direction and the effects of independence testing on smaller datasets.
@@ 3.4.6 Additive Nonlinear Time Series Causal Models
Chu et al. (Chu et al., 2008) presents a feasible procedure for learning a class of non-linear time series structures, called additive non-linear time series, and demonstrates its effectiveness in extracting detailed causal information from stationary models. The authors propose a new model specification procedure that combines statistical tests for conditional independence relations and the temporal structure of time series data. Before this, no previous methods existed for non-linear systems. Based on assumptions of stationary data, the methods discuss and extend linear methods including PC and Fast Causal Inference (FCI) with joint Normal distributions. The main algorithm used is PC, which is known to be consistent only in the absence of feedback relations and latent common causes. The paper also mentions other related algorithms such as FCI and an algorithm by Richardson and Spirtes that allow for latent variables and linear feedback relations, respectively. However, no algorithm is available that is consistent for searching linear causal models when both latent variables and feedback may be present. To extend the PC and related algorithms to a larger class of systems that includes nonlinear continuous models, a more general conditional independence test is required. Therefore the paper proposes to use a new definition of the independence test used for the additive nonlinear time series model and embed it into a causal inference algorithm. The experiments generate simulation results and show that the additive non-linear algorithm performs well for trigonometric lag models and linear lag models, but less satisfactorily for polynomial lag models. The simulation results demonstrate the effectiveness of the additive non-linear algorithm in inferring causal structure from time series data, particularly for trigonometric lag models and linear lag models.
@@ 3.4.7 Additional Methods
There are several other causal discovery methods introduced for time-series data in recent years. Temporal Causal Discovery Framework (TCDF [^4] https://github.com/M-Nauta/TCDF) (Nauta et al., 2019a) is a deep learning based framework that learns a causal graph structure in observational time-series data using attention-based convolutional neural networks. The model takes in a multivariate time-series, whereas the output is a graph comprising of nodes and the edges with lagged values. TCDF uses attention mechanism to interpret potential causal edges and Permutation Importance Validation Method (PIVM) to validate if a potential cause is a true cause. TCDF has been evaluated on financial data for stock returns and FMRI data to measure brain blood flow.
Similarly, Neural Additive Vector Autoregression Models (NAVAR [^5] https://github.com/bartbussmann/NAVAR) is a causal discovery approach for time series data that can handle highly non-linear relationships between variables. NAVAR leverages deep neural networks (DNNs) to extract the (additive) Granger causal influences from the time evolution in multivariate time series data. NAVAR assumes an additive structure, where the predictions depend linearly on independent nonlinear functions of the input variables which are modeled as nonlinear functions using the neural networks. Due to this additive structure, the causal relations can be scored and ranked. NAVAR has been evaluated on various benchmark causal discovery datasets available on the CauseMe [^6] https://causeme.uv.es/ platform for the earth sciences domain. Another deep learning based causal discovery method is Causal-HMM [^7] https://github.com/LilJing/causal_hmm (Li et al., 2021a). The method proposes to predict irreversible diseases at an early stage using time series data. It is an HMM-based hidden Markov model consisting of a subset of hidden variables, and a reformulated sequential variational autoencoder (VAE) framework to learn the proposed causal hidden Markov model. The model is applied to early prediction of peripapillary atrophy on an in-house dataset of peripapillary atrophy (PPA), and promising results are achieved on out-of-distribution test data. An ablation study shows the effectiveness of each component while visualizing the accurate identification of lesion regions.

### v2

@@ 3.3 Common Approaches
Causal discovery techniques are typically divided into three main types or approaches:
@@
This is a widely used method in causal discovery, which identifies the causal relationships by testing for conditional independence among variables. Within this framework, the Peter-Clark (PC) algorithm (Spirtes et al., 2000) is a foundational technique. It operates under the assumption of faithfulness, meaning all dependencies in a Directed Acyclic Graph (DAG) must comply with the d-separation principle. Another prominent method in this category is the Fast Causal Inference (FCI) algorithm (Spirtes et al., 2000). Unlike the PC algorithm, FCI does not assume causal sufficiency, acknowledging the possibility of unobserved confounders in the analysis of real-world data.
@@
The score-based approach is a prevalent method in causal discovery that seeks to identify the causal graph that best fits the data according to a given score function, such as the Bayesian Information Criterion (BIC). This approach requires to search across all potential graphs to maximize the score, reflecting the graph’s fit to the data (Triantafillou and Tsamardinos, 2016). Due to the nature of the approach, it can be computationally intensive, especially as the number of variables increases, leading to an exponential growth in the number of possible graphs (Nogueira et al., 2021). A notable algorithm within this category is the Greedy Equivalence Search (GES) algorithm (Chickering, 2002), which begins with an empty graph and progressively adds or removes edges based on their contribution to the overall score, optimizing the graph structure step-by-step.
@@
Functional Causal Model (FCM) approaches, also referred to as SCM-based methods, define causal relationships using specific functional forms, such as a set of equations where the variables are functions of their causal parents plus an independent noise term ($y=f(pa_{y})+e$). FCM-based techniques aim to differentiate between different DAGs within the same Markov equivalence class by making assumptions about the underlying data generation functions. A diverse array of FCM-based approaches exist for both temporal and i.i.d. data, which address linear and non-linear causal relationships through methods such as Independent Component Analysis (ICA) (Stone, 2004) and Additive Noise Models (ANMs) (Hoyer et al., 2008) respectively.
@@ 3.4 Time-Series Causal Discovery
Time-series causal discovery extends the causal structure learning process to temporal data. Causal discovery in time-series data involves dealing with the specific challenges of temporal dependencies and potential feedback loops. It has been widely used in econometrics, neuroscience, and climate science (Spirtes et al., 2000; Zhou et al., 2014; Van Nes et al., 2015; Ghysels et al., 2016). Identifying time lags is an important aspect of causal discovery in time-series data. It involves determining the delay or time difference between the cause and its effect. Time lags are crucial for understanding the dynamics of the causal relationship and for making accurate predictions in time-series forecasting tasks. To identify time lags in causal relationships, researchers may use methods like cross-correlation analysis, autocorrelation, or time-series cross-validation techniques (Spirtes et al., 2000). These methods help measure the strength of the association between variables at different time lags and assist in determining the optimal lag that maximizes the predictive performance or the strength of the causal relationship. We explain some of the widely used causal discovery methods below, along with their variants, underlying causal assumptions and limitations. A tabulated summary of these time-series causal discovery methods is given in Table 5.
@@ 3.4.1 Granger Causality, V-Granger
Granger causality (GC) is a statistical hypothesis test to determine the significance one time series in predicting another Granger (1969). It is named after Clive Granger, a Nobel laureate economist who developed the concept. A time-series X ”granger-causes” time-series Y if historic values of X play a significant role in predicting future values of Y. Granger causality is a statistical test and does not guarantee a true causal relation between two time-series. Nevertheless, Granger causality is commonly used in various fields, including economics, finance, engineering, and neuroscience, to explore relationships between time-varying data. However, it has limitations and assumptions. For instance, the choice of time lags can influence the results. GC assumes causal sufficiency condition, stationary and linear relationship and does not predict the direction of causal relation. Further, the method is data hungry and requires sufficiently large sample size.
While Granger Causality is widely used in various fields to understand the causal relationship between time series, it can yield misleading results in the presence of unobserved confounders in observational data. V-Granger (Meng, 2019), overcomes the causal sufficiency limitation by estimating Granger causality in the presence of unobserved confounders using a variational autoencoder (VAE). V-Granger incorporates an inference network to estimate the intractable posterior distribution while its generative network models the non-linear causal relationship in time-series data. Through experiments on synthetic and semi-synthetic datasets, as well as two real datasets: Butter and Temperature, the authors demonstrated how V-Granger outperforms the traditional Granger test in some cases. However, the method is not suitable for datasets with high noise levels.
@@ 3.4.2 PCMCI, PCMCI+, LPCMCI
PCMCI1 (Runge et al., 2019a) is a method for estimating causal graphs from multivariate time series datasets. In addition to causal structure learning, PCMCI can also estimate direct and mediated causal effects. A baseline of PCMCI was first introduced by (Runge et al., 2015) for identifying causal pathways in spatiotemporal climate systems. The method combines the PC stable causal discovery algorithm with momentary conditional independence (MCI) in a two-step approach. The first step is to find lagged parents of all individual nodes using the PC stable method. The second step is to test the momentary conditional independence. As a result of the second step, if two nodes are independent, then there exists no cause-effect relationship between them. The outcome of this two-step method includes a causal adjacency matrix, corresponding lag values and strength of identified relationships given by p-values. PCMCI supports multiple linear and non-parametric conditional independence tests depending on the data distribution. For instance, the partial correlation or $ParCorr()$ test assumes linear dependencies and Gaussian noise in the data. There are multiple variants of PCMCI available for different downstream tasks. For instance, PCMCI+ extends the method for auto-correlated nonlinear time series data to discovers instantaneous and lagged causal relations. For details on PCMCI’s implementation, independence tests available and different variants of PCMCI, please refer to their Github [^1] https://github.com/jakobrunge/tigramite repository.
Similarly, Latent PCMCI or LPCMCI1 (Gerhardus and Runge, 2020) is a constraint-based approach which is a variant of the PCMCI algorithm for causal discovery from observational time series data. This method is suitable when the data is non-linear, comprises hidden confounders, and the goal is to learn lagged and contemporaneous causal relations.
@@ 3.4.3 LinGAM, VARLiNGAM, TS-LiNGAM
LinGAM [^2] https://lingam.readthedocs.io/en/latest/index.html (Linear Non-Gaussian Acyclic Model) (Shimizu et al., 2006) is a framework used for causal discovery and estimation of causal effects from observational data. It focuses on identifying linear causal relationships among variables assuming non-Gaussian noise distributions and acyclic causal relations. VARLiNGAM Hyvärinen et al. (2010) extends the foundational LiNGAM model to encompass scenarios involving time series. This augmentation involves integrating the core LiNGAM model with the conventional vector autoregressive models (VAR). Consequently, it facilitates the exploration of causal relationships encompassing both delayed and concurrent (instantaneous) effects. This stands in contrast to the traditional VAR framework, which solely examines causal relations with a time delay. VARLiNGAM holds similar causal assumptions as the base LinGAM model, such as, data linearity, causal sufficiency with no hidden confounders, acyclicity of contemporaneous causal relations and non-gaussian continuous error variables.
TS-LiNGAM (Hyvärinen et al., 2008) is an extension of the LiNGAM algorithm for estimating both instantaneous and lagged causal edges in time-series data. The model is a combination of autoregressive models and structural equation modeling (SEM). It assumes the external noises to be mutually independent, temporally uncorrelated, and to be non-Gaussian. The overall method has four steps: i) Estimation of a classic autoregressive model, ii) Residuals computation, iii) Performing LiNGAM analysis on the residuals, and iv) Computation of the causal effect estimates. The method has been evaluated on real datasets such as financial (stock) data and magnetoencephalographic data.
@@ 3.4.4 TiMiNo
TiMINo (Peters et al., 2013) approach provides a new method at that time for causal inference on time series data that overcomes some of the methodological issues of previous methods that could not capture nonlinear and instantaneous effects. Specifically, TiMINo extends Structural Equation Models (SEMs) for time series analysis, utilizing a restricted function class to achieve general identifiability results. This approach accounts for both lagged and instantaneous effects, which can be nonlinear and unfaithful, as well as non-instantaneous feedback within the time series. Given a sample of finite and continuous-distributed multivariate time series data, and output the summary time graph if the input data satisfies model assumptions. If the input does not meet the assumptions, TiMINo typically refrains from making causal determinations, avoiding incorrect causal conclusions. The model is represented as $X_{t}^{i}=f_{i}((\mathbf{PA_{p}^{i}})_{t-p},\ldots,(\mathbf{PA_{1}^{i}})_{t-1},(\mathbf{PA_{0}^{i}})_{t},N_{t}^{i})$, where $p$ is the time lag for independence test, $(\mathbf{PA_{1}^{i}})$ is the set of parent nodes for node $i$, $X_{t}$ is time-series data with finite and continuous distribution, and $N_{t}^{i}$ is a noise vector jointly independent over i and t. The prior assumption of this approach is that each time sequence can be represented by a function of all its direct causes and some noise variable, and restricts the function class to provide general identifiability results. Potential weaknesses of TiMINo are discussed including the possibility of fitting a model in the wrong direction and the effects of independence testing on smaller datasets.
@@ 3.4.6 Additive Nonlinear Time Series Causal Models
Chu et al. (Chu et al., 2008) introduced a method for learning additive non-linear time series, and demonstrated its effectiveness in extracting detailed causal information from stationary models. Before this, no previous methods existed for non-linear systems. Based on assumptions of stationary data, the methods discuss and extend linear methods including PC and Fast Causal Inference (FCI) with joint Normal distributions. The primary algorithm employed is PC, which is known to be consistent only when there are no feedback relations or latent common causes. The paper also mentions other related algorithms such as FCI and an algorithm by Richardson and Spirtes that allow for latent variables and linear feedback relations, respectively. However, no algorithm consistently identifies linear causal models in the presence of latent variables and feedback. To extend the PC and similar algorithms to a broader range of systems, including nonlinear continuous models, a more comprehensive conditional independence test is necessary. Therefore the paper proposes to use a new definition of the independence test used for the additive nonlinear time series model and embed it into a causal inference algorithm. The simulation results demonstrate the effectiveness of the additive non-linear algorithm in inferring causal structure from time series data, particularly for trigonometric lag models and linear lag models.
@@ 3.4.7 Additional Methods
There are several other causal discovery methods introduced for time-series data in recent years. Temporal Causal Discovery Framework (TCDF [^4] https://github.com/M-Nauta/TCDF) (Nauta et al., 2019a) aims to uncover causal graph structures within observational time-series data by utilizing attention-based convolutional neural networks. This model processes multivariate time-series data and outputs a graph that includes nodes and edges with corresponding lagged values. TCDF employs the Permutation Importance Validation Method (PIVM) to validate potential causes. In the paper, TCDF’s performance has been evaluated using financial stock return data and fMRI data for tracking brain blood flow.
Similarly, Neural Additive Vector Autoregression Models (NAVAR [^5] https://github.com/bartbussmann/NAVAR) discover highly non-linear causal relationships between factors for time series data. NAVAR employs deep neural networks to derive the additive Granger causal influences from the temporal dynamics of multivariate time series data. As the model name implies, NAVAR proposes an additive structure to extend VAR. The new method makes predictions linearly depend on independent nonlinear functions of the input variables, modeled through a multi-layer perceptron. This approach enables the scoring and ranking of causal relationships. NAVAR has been validated using various benchmark datasets for causal discovery available on the CauseMe [^6] https://causeme.uv.es/ platform, particularly in the fields of Earth science.
TS-CausalNN (Faruque et al., 2024) learns causal graphs with lagged and instantaneous causal links simultaneously from non-stationary non-linear multivariate time series data. TS-CausalNN utilizes the concept of learning non-stationary features using non-linear functions and proposes a novel parallel neural network design using a custom causal convolution 2D layer. The model processes temporal data without any external adjustments like detrending, deseasonalizing, etc., and generates a full causal graph with link values. Three synthetic datasets and three real-world datasets with non-stationary and non-linear features were used to assess the effectiveness of this approach.
Another causal discovery method based on deep learning is Causal-HMM [^7] https://github.com/LilJing/causal_hmm (Li et al., 2021a). The method proposes to predict irreversible diseases at an early stage using time series data. It is an HMM-based hidden Markov model consisting of a subset of hidden variables, and a reformulated sequential variational autoencoder (VAE) framework to learn the proposed causal hidden Markov model. The model is applied to early prediction of peripapillary atrophy on an in-house dataset of peripapillary atrophy (PPA).